# OmniHands: Towards Robust 4D Hand Mesh Recovery via A Versatile Transformer

### v3

@@
Abstract
In this paper, we introduce OmniHands, a universal approach to recovering interactive hand meshes and their relative movement from monocular or multi-view inputs. Our approach addresses two major limitations of previous methods: lacking a unified solution for handling various hand image inputs and neglecting the positional relationship of two hands within images. To overcome these challenges, we develop a universal architecture with novel tokenization and contextual feature fusion strategies, capable of adapting to a variety of tasks. Specifically, we propose a Relation-aware Two-Hand Tokenization (RAT) method to embed positional relation information into the hand tokens. In this way, our network can handle both single-hand and two-hand inputs and explicitly leverage relative hand positions, facilitating the reconstruction of intricate hand interactions in real-world scenarios. As such tokenization indicates the relative relationship of two hands, it also supports more effective feature fusion. To this end, we further develop a 4D Interaction Reasoning (FIR) module to fuse hand tokens in 4D with attention and decode them into 3D hand meshes and relative temporal movements. The efficacy of our approach is validated on several benchmark datasets. The results on in-the-wild videos and real-world scenarios demonstrate the superior performances of our approach for interactive hand reconstruction. More video results can be found on the project page: https://OmniHand.github.io.
Figure 1: The proposed method, OmniHands, can robustly recover interactive hand meshes and their relative movement from monocular inputs.
11footnotetext: Corresponding authors.
1 Introduction
@@
Despite the remarkable progress achieved by existing methods [32, 34, 64] in this field, three major challenges remain unresolved. First, there is a lack of a universal method that can handle various input statuses for both single-handed and two-handed input. When using existing single-hand reconstruction models to reconstruct two hands, while they can individually reconstruct each hand, they fail to address the issue of hand occlusion and cannot compute the relative distance between the hands. On the other hand, current two-hand reconstruction models, due to their use of single images containing both hands as input, are limited to reconstructing scenarios where the hands are in close proximity, thereby exhibiting poor generalizability. Second, current hand recovery methods address specific subproblems such as monocular hand reconstruction, video-based reconstruction, and multi-view reconstruction, however, a unified and general model to handle these tasks is still lacking. Third, although current hand reconstruction methods perform well on test data, their effectiveness in in-the-wild scenarios is limited due to the inadequate size of training datasets, resulting in poor practical value.
The success of recent approaches in transformer-based large-scale models such as GPT-4[1], ViTpose [61], HaMeR [45] provided us with the insight to investigate the usage of simple and high-capacity models along with large-scale in-the-wild data. Therefore, we proposed OmniHands, a transformer-based universal solution to recover interactive hand meshes cross omniform scenarios. The transformer possesses versatility, enabling it to perform unified forward operations on different input forms such as single-frame input, temporal input, and multi-view input. Additionally, multi-layer transformers’ ability to process large datasets allows us to train a generalized model that remains robust in real-world scenarios across various lighting conditions, backgrounds, and interaction poses.
@@
  • Our final solution, OmniHands, achieves robust hand mesh recovery in real-world scenarios, serving as a versatile solution to handle hand inputs in various forms, whether the images contain single or two hands, in single frames or temporal sequences or multi-view sequences.
2 Related Work
2.1 Monocular Hand Mesh Recovery
@@
Recent years have witnessed great progress in hand mesh recovery from monocular RGB input. Following the advances in monocular human mesh recovery [25, 29, 28, 67, 66, 52], some researchers [2, 45, 42, 70, 3] use neural networks to regress the driving parameters of hand templates such as MANO [48] to recover hand meshes, while others prefer to directly estimate the vertex coordinates of hand meshes through GCN [18, 30, 51, 7, 5, 6] or transformer [35, 34].
@@
Recently, there have also been some efforts to simultaneously recover both two hands of humans. However, two hands mesh recovery is much more challenging than one hand, due to larger pose distribution and more severe occlusion. Furthermore, compared to the large-scale single-hand datasets [45] with rich capture scenarios and postures, existing two-hand datasets are much fewer, and most of them are captured in limited environments.
@@
A straightforward method for two-hand mesh recovery is to locate and crop each hand separately from the input image, and then convert the task into single-hand scenarios by processing two cropped hand images individually. This strategy is wildly used in full-body pose estimation frameworks [68, 24, 9, 49, 15, 40, 71]. However, this approach is still based on single-hand recovery network, which usually performs poorly under close interaction situations. Moreover, the relative translation between two hands is ambiguous when the two hands are processed separately.
2.2 Hand Mesh Recovery under Interaction
@@
To better address two-hands pose estimation and mesh recovery in close interaction scenarios, some researchers [32, 65, 64, 76, 26, 39, 37, 33] used one image containing both hands as input. Zhang et al. [65] utilized 2d joints projection attention and context-aware refinement module to improve the pose accuracy. Li et al. [32] utilize transformer module [12] to learn implicit attention between two hands. Zuo et al. [76] propose a data-based prior of two interacting hands. Yu et al. [64] constructed the attention relationship between the two hands based on 2D hand center maps. Although the above approaches have produced promising results, there are still some limitations. Because these methods use a single image containing both hands as input, when the two hands are far apart, the area proportion of hands in the input image will be too small, which can reduce the accuracy of mesh recovery. Furthermore, their network design inherently lacks the ability to recover single-hand mesh and only can operate under two-hand scenarios.
@@
Therefore, we propose a new network that is compatible with both single-hand recovery and two-hand recovery. We take one or two cropped hand images as input, enabling utilize large-scale single-hand datasets to boost two-hand recovery results. We also tokenize the 2D relative distance and overlap between two cropped hand images by a powerful transformer [56, 12] module, resulting in accurate relative root translation estimation and interacting hands mesh recovery.
2.3 Temporal Hand Mesh Recovery
@@
Several methods have been proposed for hand mesh recovery from temporal sequence data [36, 22, 62, 16, 27, 8]. Hasson et al. [22] leverage the optical flow of nearby frames to constrain the motion of hands and objects by photometric supervision. SeqHAND [62] incorporates the convolution-LSTM layer to capture the sequential relationship between consecutive hand poses. Deformer [16] proposes a novel dynamic fusion module that explicitly deforms nearby frames with clear hand visibility for robust hand pose estimation from temporal input. VIBE [27] and TCMR [8] utilize the recurrent neural network to produce temporally consistent motion.
@@
2.4 Multi-View Hand Mesh Recovery
Recovering hands from multi-view input is a similar task as multi-view human reconstruction. Many recent researches have been proposed for multi-view 3D human pose estimation  [46, 11, 53, 44, 72]. Some of the methods [46, 11, 53] focus on matching the projection of 3D key-points with their corresponding 2D key-points from multiple camera views, and some [44, 72] focus on fusing the 2D visual features to build a 3D structure.
Recent research on multi-view hand reconstruction has seen various innovative approaches. Some researchers [10, 20] utilize implicit fields to recover the shape and appearance of hands. Yang et al. [63] employ hand vertices as a point cloud embedded in the intersection volume of frustums from multiple cameras to integrate multi-view information. Wang et al. [58] adapt multi-view hand data to enhance the performance of single-view hand reconstruction. Additionally, Gan et al. [17] propose a neural rendering scheme to refine both the hand mesh and textures. However, most multi-view methods require extrinsic and intrinsic calibration information for each camera to accurately determine the 3D positions of hands. In contrast, our approach is capable of utilizing only visual information to recover precise hand meshes.
3 Method
Figure 2: Overview of our OmniHands framework. OmniHands is a transformer-based network, which takes various forms of inputs and estimates two-hand meshes with their relative positions. General modules are used to process different forms of input data. For simplicity of representation, since the hand tokens of both hands will undergo the same processing in FIR, the hand tokens $G^{*},L^{*}$ in the diagram represent those of one hand. In the top left corner we present our training set, indicating the data volume and whether including Interaction, Time sequences, Multi-view, Real-World data.
@@ 3 Method
3.1 Problem Formulation
Given a single-frame or a temporal sequence or a multi-view sequence of images $I$, our goal is to reconstruct the hand meshes and their relative position. We leverage the widely used parametric hand model MANO [48]. MANO defines a function $\mathbf{Mn}(\theta,\beta)$ that takes two parameters pose $\theta\in\mathbb{R}^{48}$ and shape $\beta\in\mathbb{R}^{10}$ as input, returns 3D hand mesh $V\in\mathbb{R}^{778\times 3}$ and 3D joint keypoints $V\in\mathbb{R}^{21\times 3}$. The outputs of our model are MANO parameters of two hands: $\theta_{R},\theta_{L},\beta_{R},\beta_{L}$, by which we can regress meshes and joints of each hand, and a vector $\upsilon$ to represent the joint root distance of two hands. We define $\upsilon$ to be the location of the left root joint location when we subtract the joints of both hands from the right hand’s 3D root joint location. Note that our method can robustly handle both single-hand and two-hand inputs. When there is only a single hand visible in the images, our method is simply degraded to the estimation of the mesh of one hand.
3.2 Relation-aware Two-Hand Tokenization
Existing approaches to hand mesh recovery typically take the cropped hand images as input, but neglect the hand relationship in the images. Besides, these approaches handle two-hand and single-hand scenarios in different manners, resulting in jittering artifacts when the two hands operate in interaction and separately. To handle these issues in our transformer-based architecture, we propose a Relation-aware Two-Hand Tokenization (RAT) method to take the relation of two hands into consideration and handle inputs in various scenarios. Following recent state-of-the-art practices in 3D human/hand mesh recovery [60, 19, 45], we utilize ViT [12] as the visual encoder to tokenize image patches. In contrast to these methods, we handle both the single-hand and two-hand inputs in a unified manner and consider embedding positional relation information of each hand into its tokens during the tokenization process.
@@ 3.2 Relation-aware Two-Hand Tokenization
Specifically, for each input image, we follow ViT [12] to segment it into $H\times W$ image patches and subsequently feed them into the ViT backbone, resulting in a vision token map $M\in\mathbb{R}^{H\times W\times C}$ To enrich each token map with relationship information, we further calculate relation maps $\{\mathcal{R}_{R2L},\mathcal{R}_{L2R}\}$ and embed them into the final image tokens, so that they can contain both the patch-wise visual knowledge and patch-to-patch relation information. Basically, the relation maps are the composition of the relative distance map and the overlapping map. Without losing generalization, we describe the right-hand relation map $\mathcal{R}_{R2L}$ in detail as follows. The left-hand relation map $\mathcal{R}_{L2R}$ can be calculated similarly.
@@
The relative distance map is calculated based on the hand position in the original images. Let $\mathcal{B}=[c_{x},c_{y},s_{x},s_{y}]$ be the bounding box of one hand, $c_{x},c_{y}$ the x-coordinate and y-coordinate of the center position, and $s_{x},s_{y}$ the box scale lengths on x-axis and y-axis. We obtain a position map $\mathcal{C}\in\mathbb{R}^{H\times W\times 2}$, representing positions of the center in each image patch. For the patch $p(i,k)$ at the $i$-th column and the $k$-th row, its value is
$\mathcal{C}(p)=[c_{x}+\frac{(2i-W)\cdot s_{x}}{2W},c_{y}+\frac{(2k-H)\cdot s_{y}}{2H}].$ (1)
@@
$\mathcal{D}_{R2L}(p)=\psi(\tau\cdot\frac{\mathcal{C}_{R}(p)-\mathcal{C}_{L}(p)}{s})$ (2)
@@
$\mathcal{O}_{R2L}(p)=\begin{cases}1,&Inside(p,\mathcal{B}_{L})\\ -1,&\text{otherwise}\end{cases}$ (3)
@@
$\mathcal{R}_{R2L}=MLP(\mathcal{D}_{R2L}\oplus\mathcal{D}_{L2R}\oplus\mathcal{O}_{R2L}),$ (4)
@@
$\mathcal{M_{R}}=Flatten(M_{R}+\mathcal{R}_{R2L})$ (5)
@@
$\mathcal{M^{*}}=FFN(Softmax(\frac{Q(\mathcal{M})K(\mathcal{M})}{\sqrt{dim(\mathcal{M})})})V(\mathcal{M}))$ (6)
@@
$G=FFN(Softmax(\frac{Q(\alpha)K(\mathcal{M}^{*})}{\sqrt{dim(\mathcal{M}^{*})})})V(\mathcal{M}^{*}))$ (7)
Then the hand tokens are taken for further feature fusion according to the specific task demands.
3.3 4D Interaction Reasoning
@@
$G^{*}=FFN(Softmax(\frac{Q(\alpha)K(\mathcal{G})}{\sqrt{dim(\mathcal{G})})})V(\mathcal{G})$ (8)
@@
$\mathcal{L}^{*}=FFN(Softmax(\frac{Q(\mathcal{G})K(\mathcal{G})}{\sqrt{dim(\mathcal{G})})})V(\mathcal{G})$ (9)
@@
Although we predict the root orientations and offset vectors for each frame individually, the relative orientation and distance between the right hand and left hand should remain invariant across all frames. This means that when we transform all right hands to the zero position and zero rotation, the positions and orientations of the left hands should remain consistent. Therefore, we apply a constraint (Eq 17) to maintain this invariance.
3.4 Losses

### v4

# OmniHands: Robust Motion Capture of Interactive Hands via A Versatile Transformer

@@
Abstract.
In this paper, we introduce OmniHands, a universal approach to recovering interactive hand motions and their relative movement from monocular or multi-view inputs. Our approach addresses three major limitations of previous methods: lacking a unified solution for handling single hand and interactive hands image inputs, lacking a versatile framework to support various 4D tasks such as video-based tasks and multi-view tasks, and unable to maintain robust performance in real-world data. To overcome these challenges, we develop a universal architecture with novel tokenization and contextual feature fusion strategies, capable of adapting to a variety of tasks. Specifically, we propose a Relation-aware Two-Hand Tokenization (RAT) method to embed positional relation information into the hand tokens. In this way, our network can handle both single-hand and two-hand inputs and explicitly leverage relative hand positions, facilitating the reconstruction of intricate hand interactions in real-world scenarios. As such tokenization indicates the relative relationship of two hands, it also supports more effective feature fusion. To this end, we further develop a 4D Interaction Reasoning (FIR) module to fuse hand tokens in 4D with attention and decode them into 3D hand meshes and relative temporal movements. The efficacy of our approach is validated on several benchmark datasets. The results on in-the-wild videos and real-world scenarios demonstrate the superior performances of our approach for interactive hand reconstruction. The code and video results can be found in omnihands.github.io.
Hand motion capture, two-hand interaction, multi-view, temporal
journal: TOGjournalvolume: 45journalnumber: 4article: 39copyright: acmlicensedccs: Computing methodologies Motion capture
Figure 1. The proposed method, OmniHands, can robustly recover interactive hand motions and their relative movement from monocular inputs.
1. Introduction
@@
Despite the remarkable progress achieved by existing methods (Li et al., 2022; lin2021metro; Yu et al., 2023) in this field, three major challenges remain unresolved. First, there is a lack of a universal method that can handle various input statuses for both single-handed and two-handed input. When using existing single-hand reconstruction models to reconstruct two hands, while they can individually reconstruct each hand, they fail to address the issue of hand occlusion and cannot compute the relative distance between the hands. On the other hand, current two-hand reconstruction models, due to their use of single images containing both hands as input, are limited to reconstructing scenarios where the hands are in close proximity, thereby exhibiting poor generalizability. Second, current hand recovery methods address specific subproblems such as monocular hand reconstruction, video-based reconstruction, and multi-view reconstruction, however, a unified and general model to handle these tasks is still lacking. Third, although current hand reconstruction methods perform well on test data, their effectiveness in in-the-wild scenarios is limited due to the inadequate size of training datasets, resulting in poor practical value.
The success of recent approaches in transformer-based large-scale models such as GPT-4(achiam2023gpt), ViTpose (Xu et al., 2022), HaMeR (Pavlakos et al., 2023) provided us with the insight to investigate the usage of simple and high-capacity models along with large-scale in-the-wild data. Therefore, we proposed OmniHands, a transformer-based universal solution to recover interactive hand meshes cross various scenarios. The transformer possesses versatility, enabling it to perform unified forward operations on different input forms such as single-frame input, temporal input, and multi-view input. Additionally, multi-layer transformers’ ability to process large datasets allows us to train a generalized model that remains robust in real-world scenarios across various lighting conditions, backgrounds, and interaction poses.
@@
  • Our final solution, OmniHands, achieves robust hand motion capture in real-world scenarios, serving as a versatile solution to handle hand inputs in various forms, whether the images contain single or two hands, in single frames or temporal sequences or multi-view sequences.
2. Related Work
2.1. Monocular Hand Motion Capture
@@
Recent years have witnessed great progress in hand mesh recovery from monocular RGB input. Following the advances in monocular human mesh recovery (kanazawa2018end; kolotouros2019graphcmr; kolotouros2019learning; zhang2020densepose2smpl; zhang2021pymaf; tian2023recovering), some researchers (Baek et al., 2019; Pavlakos et al., 2023; Park et al., 2022; Zhang_2019_ICCV; boukhayma20193d) use neural networks to regress the driving parameters of hand templates such as MANO (MANO:SIGGRAPHASIA:2017) to recover hand meshes, while others prefer to directly estimate the vertex coordinates of hand meshes through GCN (Ge_2019_CVPR; 2020Weakly; tang2021towards; Choi et al., 2020; Chen et al., 2021, 2022) or transformer (Lin et al., 2021; lin2021metro).
@@
Recently, there have also been some efforts to simultaneously recover both two hands of humans. However, two hands mesh recovery is much more challenging than one hand, due to larger pose distribution and more severe occlusion. Furthermore, compared to the large-scale single-hand datasets (Pavlakos et al., 2023) with rich capture scenarios and postures, existing two-hand datasets are much fewer, and most of them are captured in limited environments.
@@
A straightforward method for two-hand mesh recovery is to locate and crop each hand separately from the input image, and then convert the task into single-hand scenarios by processing two cropped hand images individually. This strategy is widely used in full-body pose estimation frameworks (pymafx2023; Joo_2018_CVPR; Choutas et al., 2020; rong2020frankmocap; Feng et al., 2021; Moon et al., 2022; lightcap2021). However, this approach is still based on single-hand recovery network, which usually performs poorly under close interaction situations. Moreover, the relative translation between two hands is ambiguous when the two hands are processed separately.
2.2. Hand Motion Capture under Interaction
@@
To better address two-hands pose estimation and mesh recovery in close interaction scenarios, some researchers (Li et al., 2022; Zhang et al., 2021; Yu et al., 2023; Zuo et al., 2023; kim2021end; Moon_2020_ECCV_InterHand2.6M; meng2022hdr; Li et al., 2024) used one image containing both hands as input. Zhang et al. (Zhang et al., 2021) utilized 2d joints projection attention and context-aware refinement module to improve the pose accuracy. Li et al. (Li et al., 2022), Cho et al. (cho2023transformer) and Jiang et al. (jiang2023a2j) utilize transformer module (Dosovitskiy et al., 2020) to learn implicit attention between two hands. Zuo et al. (Zuo et al., 2023) propose a data-based prior of two interacting hands. Yu et al. (Yu et al., 2023) constructed the attention relationship between the two hands based on 2D hand center maps. Although the above approaches have produced promising results, there are still some limitations. Because these methods use a single image containing both hands as input, when the two hands are far apart, the area proportion of hands in the input image will be too small, which can reduce the accuracy of mesh recovery. Moon et al. (Moon, 2023) separates the two hands as input, but its prediction of the relative position between the hands relies on bounding box information rather than visual cues from the image. Moreover, it also requires input images that contain both hands since the separation is conducted during forward process. Furthermore, these above network designs inherently lack the ability to recover single-hand mesh and only can operate under two-hand scenarios.
@@
Therefore, we propose a new network that is compatible with both single-hand recovery and two-hand recovery. We take one or two cropped hand images as input, enabling utilize large-scale single-hand datasets to boost two-hand recovery results. We also tokenize the 2D relative distance and overlap between two cropped hand images by a powerful transformer (Vaswani et al., 2017; Dosovitskiy et al., 2020) module, resulting in accurate relative root translation estimation and interacting hands mesh recovery.
2.3. Temporal Hand Motion Capture
@@
Several methods have been proposed for hand mesh recovery from temporal sequence data (Liu et al., 2021; Hasson et al., 2020; Yang et al., 2020; Fu et al., 2023; Kocabas et al., 2020; Choi et al., 2021). RGB2Hands (wang2020rgb2hands) proposed a real-time hand tracking approach that utilizes both segmentation and depth prediction to enhance performance. Hasson et al. (Hasson et al., 2020) leverage the optical flow of nearby frames to constrain the motion of hands and objects by photometric supervision. SeqHAND (Yang et al., 2020) incorporates the convolution-LSTM layer to capture the sequential relationship between consecutive hand poses. Deformer (Fu et al., 2023) proposes a novel dynamic fusion module that explicitly deforms nearby frames with clear hand visibility for robust hand pose estimation from temporal input. VIBE (Kocabas et al., 2020) and TCMR (Choi et al., 2021) utilize the recurrent neural network to produce temporally consistent motion.
@@
2.4. Multi-View Hand Motion Capture
Recovering hands from multi-view input is a similar task as multi-view human reconstruction. Many recent researches have been proposed for multi-view 3D human pose estimation  (qiu2019cross; dong2019fast; tu2020voxelpose; pavlakos2017harvesting; zhang2021adafuse). Some of the methods (qiu2019cross; dong2019fast; tu2020voxelpose) focus on matching the projection of 3D key-points with their corresponding 2D key-points from multiple camera views, and some (pavlakos2017harvesting; zhang2021adafuse) focus on fusing the 2D visual features to build a 3D structure.
Recent research on multi-view hand reconstruction has seen various innovative approaches. Some researchers (corona2022lisa; guo2023handnerf) utilize implicit fields to recover the shape and appearance of hands. Yang et al.(yang2023poem) employ hand vertices as a point cloud embedded in the intersection volume of frustums from multiple cameras to integrate multi-view information. Wang et al.(wang2024sima) adapt multi-view hand data to enhance the performance of single-view hand reconstruction. Additionally, Gan et al.(gan2024fine) propose a neural rendering scheme to refine both the hand mesh and textures. However, most multi-view methods require extrinsic and intrinsic calibration information for each camera to accurately determine the 3D positions of hands. In contrast, our approach is capable of utilizing only visual information to recover precise hand meshes.
3. Method
Figure 2. Overview of our OmniHands framework. OmniHands is a transformer-based network, which takes various forms of inputs and estimates two-hand meshes with their relative positions. General modules are used to process different forms of input data. For simplicity of representation, since the hand tokens of both hands will undergo the same processing in FIR, the hand tokens $G^{*},L^{*}$ in the diagram represent those of one hand. In the top left corner we present our training set, indicating the data volume and whether including Interaction, Time sequences, Multi-view, Real-World data.
A Figure.
@@ 3. Method
3.1. Problem Formulation
Given a single-frame or a temporal sequence or a multi-view sequence of images $I$, our goal is to reconstruct the hand meshes and their relative position. We leverage the widely used parametric hand model MANO (MANO:SIGGRAPHASIA:2017). MANO defines a function $\mathbf{Mn}(\theta,\beta)$ that takes two parameters pose $\theta\in\mathbb{R}^{48}$ and shape $\beta\in\mathbb{R}^{10}$ as input, returns 3D hand mesh $V\in\mathbb{R}^{778\times 3}$ and 3D joint keypoints $V\in\mathbb{R}^{21\times 3}$. The outputs of our model are MANO parameters of two hands: $\theta_{R},\theta_{L},\beta_{R},\beta_{L}$, by which we can regress meshes and joints of each hand, and a vector $\upsilon$ to represent the joint root distance of two hands. We define $\upsilon$ to be the location of the left root joint location when we subtract the joints of both hands from the right hand’s 3D root joint location. Note that our method can robustly handle both single-hand and two-hand inputs. When there is only a single hand visible in the images, our method is simply degraded to the estimation of the mesh of one hand.
3.2. Relation-aware Two-Hand Tokenization
Existing approaches to hand mesh recovery typically take the cropped hand images as input, but neglect the hand relationship in the images. Besides, these approaches handle two-hand and single-hand scenarios in different manners, resulting in jittering artifacts when the two hands operate in interaction and separately. To handle these issues in our transformer-based architecture, we propose a Relation-aware Two-Hand Tokenization (RAT) method to take the relation of two hands into consideration and handle inputs in various scenarios. Following recent state-of-the-art practices in 3D human/hand mesh recovery (Xu et al., 2022; Goel et al., 2023; Pavlakos et al., 2023), we utilize ViT (Dosovitskiy et al., 2020) as the visual encoder to tokenize image patches. In contrast to these methods, we handle both the single-hand and two-hand inputs in a unified manner and consider embedding positional relation information of each hand into its tokens during the tokenization process.
@@ 3.2. Relation-aware Two-Hand Tokenization
Specifically, for each input image, we follow ViT (Dosovitskiy et al., 2020) to segment it into $H\times W$ image patches and subsequently feed them into the ViT backbone, resulting in a vision token map $M\in\mathbb{R}^{H\times W\times C}$ . To enrich each token map with relationship information, we further calculate relation maps $\{\mathcal{R}_{R2L},\mathcal{R}_{L2R}\}$ and embed them into the final image tokens, so that they can contain both the patch-wise visual knowledge and patch-to-patch relation information. Basically, the relation maps are the composition of the relative distance map and the overlapping map. Without losing generalization, we describe the right-hand relation map $\mathcal{R}_{R2L}$ in detail as follows. The left-hand relation map $\mathcal{R}_{L2R}$ can be calculated similarly.
@@
The relative distance map is calculated based on the hand position in the original images.
Let $\mathcal{B}=[c_{x},c_{y},s_{x},s_{y}]$ be the bounding box of one hand, $c_{x},c_{y}$ the x-coordinate and y-coordinate of the center position, and $s_{x},s_{y}$ the box scale lengths on x-axis and y-axis. We obtain a position map $\mathcal{C}\in\mathbb{R}^{H\times W\times 2}$, representing positions of the center in each image patch. For the patch $p(i,k)$ at the $i$-th column and the $k$-th row, its value is
(1) $\mathcal{C}(p)=[c_{x}+\frac{(2i-W)\cdot s_{x}}{2W},c_{y}+\frac{(2k-H)\cdot s_{y}}{2H}].$
@@
(2) $\mathcal{D}_{R2L}(p)=\psi(\tau\cdot\frac{\mathcal{C}_{R}(p)-\mathcal{C}_{L}(p)}{s})$
@@
(3) $\mathcal{O}_{R2L}(p)=\begin{cases}1,&Inside(p,\mathcal{B}_{L})\\ -1,&\text{otherwise}\end{cases}$
@@
(4) $\mathcal{R}_{R2L}=MLP(\mathcal{D}_{R2L}\oplus\mathcal{D}_{L2R}\oplus\mathcal{O}_{R2L}),$
@@
(5) $\mathcal{M_{R}}=Flatten(M_{R}+\mathcal{R}_{R2L})$
@@
(6) $\mathcal{M^{*}}=FFN(Softmax(\frac{Q(\mathcal{M})K(\mathcal{M})}{\sqrt{dim(\mathcal{M})})})V(\mathcal{M}))$
@@
(7) $G=FFN(Softmax(\frac{Q(\alpha)K(\mathcal{M}^{*})}{\sqrt{dim(\mathcal{M}^{*})})})V(\mathcal{M}^{*}))$
The vector $\alpha$ is predefined and has different value for left and right hands. Then the hand tokens are taken for further feature fusion according to the specific task demands.
3.3. 4D Interaction Reasoning
@@
(8) $G^{*}=FFN(Softmax(\frac{Q(\alpha)K(\mathcal{G})}{\sqrt{dim(\mathcal{G})})})V(\mathcal{G})$
@@
(9) $\mathcal{L}^{*}=FFN(Softmax(\frac{Q(\mathcal{G})K(\mathcal{G})}{\sqrt{dim(\mathcal{G})})})V(\mathcal{G})$
@@
Although we predict the root orientations and offset vectors for each frame individually, the relative orientation and distance between the right hand and left hand should remain invariant across all frames. This means that when we transform all right hands to the zero position and zero rotation, the positions and orientations of the left hands should remain consistent. Therefore, we apply a constraint (Eq 17) to maintain this invariance. During inference, we average the relative translations and relative orientations to ensure consistency across views.
3.4. Losses